# Floquet-Lindblad Master Equation Approach to Open Quantum System Dynamics

### v1

@@ IV Discussion
In this paper, a new type of ODE solver based on Floquet theory is discussed. A short derivation is presented, in which both a secular and nonsecular master equation of the Floquet-Lindblad type are derived. The secular version is easier to obtain and cheaper to solve computationally, but there are many cases where a relaxed secular approximation is necessary. After discussing the theory behind the FLiME, two exemplar systems were discussed. While these systems are themselves rather simple, they show some of the utility of this new solver. The first system was that of a driven 2-level system, however, instead of making the usual rotating wave approximation, the full Hamiltonian was used due to high driving power. FLiMESolve was able to address the computational overhead required of this time-dependent Hamiltonian via its efficiency and its applicability to problems in the Floquet regime of ordinary differential equations. The second system was that of the pulsed 2-level system. Since the pulses are made up of a sum of waves of different harmonic frequencies, the full secular approximation realizes a state evolution that is nonphysical because it ignores frequencies above the fundamental. However, the ability of FLiMESolve to relax the secular approximation allows for the recovery of accurate system evolution.
There are two avenues for future work on FLiMESolve. Firstly, the decay rate can be made non-constant over all frequencies. This would bring the form of the decay rate, $\gamma$, to match non-secular master equations derived elsewhere, and would allow for the simulation of even more richly dynamic systems [14]. The second avenue for future work is to incorporate multimode Floquet theory into the FLiME [25, 26]. In semiclassical multimode Floquet theory, commensurate driving frequencies (i.e. those whose quotient is a rational number) are themselves treated as the Fourier expansion of some lower frequency driving term, given by the lowest common denominator of the two (or more) commensurate frequencies. This further complicates the indexing of any eventual master equations, but would allow for the simulation of systems with many commensurate driving terms.

### v2

@@ IV Discussion
In this paper, a new type of ODE solver based on Floquet theory is discussed. A short derivation was presented, in which both a secular and nonsecular master equation of the Floquet-Lindblad type are derived. The secular version is easier to obtain and cheaper to solve computationally, but there are cases where a relaxed secular approximation is necessary. After discussing the theory behind the FLiME, two exemplar systems were discussed. While these systems are themselves rather simple, they show the utility of this new solver. The first system was that of a driven 2-level system, however, instead of making the usual rotating wave approximation, the full Hamiltonian was used due to high driving power. FLiMESolve was able to address the computational overhead required of this time-dependent Hamiltonian via its efficiency and its applicability to problems in the Floquet regime of ordinary differential equations. The second system was that of the 2-level system excited by a pulse train. Since the pulses are made up of a sum of waves of different harmonic frequencies, the full secular approximation realizes a state evolution that is nonphysical because it ignores frequencies above the fundamental. However, the ability of FLiMESolve to relax the secular approximation allows for the recovery of accurate system evolution.
There are two avenues for future work on FLiMESolve. Firstly, the decay rate can be made non-constant over all frequencies. This would bring the form of the decay rate, $\gamma$, to match non-secular master equations derived elsewhere, and would allow for the simulation of even more richly dynamic systems [15]. The second avenue for future work is to incorporate multimode Floquet theory into the FLiME [28, 29]. In semiclassical multimode Floquet theory, commensurate driving frequencies (i.e,. those whose quotient is a rational number) are themselves treated as the Fourier expansion of some lower frequency driving term, given by the lowest common denominator of the two (or more) commensurate frequencies. This further complicates the indexing of any eventual master equations, but would allow for the simulation of systems with many commensurate driving terms.